OpenStudy (xhawaiiangirlx):

Freddy has a shelf that is 72 over 5 inches wide. How many catalogs can Freddy arrange on the shelf if each catalog is 4 over 5 of an inch thick? 11, because the number of catalogs is 72 over 5 divided by 5 over 4 12, because the number of catalogs is 72 over 5 multiplied by 4 over 5 16, because the number of catalogs is 72 over 5 multiplied by 5 over 4 18, because the number of catalogs is 72 over 5 divided by 4 over 5 @the_ocean_girl @Frozen15 @mathmale

OpenStudy (theraggedydoctor):

What do you think? He has one that's \(\frac{ 72 }{ 5 }\) and you want to know how many catalogs he can fit on it if each of them are \(\frac{4}{5}\) of an inch thick.

OpenStudy (xhawaiiangirlx):

Idk im really bad at math

OpenStudy (theraggedydoctor):

Okay, well give it a shot, at least. You want to fit the smaller number onto the bigger one.. What do you think you should do? Divide or multiply?

OpenStudy (xhawaiiangirlx):

divide?

OpenStudy (theraggedydoctor):

Yes. So now we know it's either the first or last choice.

OpenStudy (xhawaiiangirlx):

last one

OpenStudy (xhawaiiangirlx):

because 72/5 divided by 4/5=18

OpenStudy (xhawaiiangirlx):

thanks for help!

OpenStudy (theraggedydoctor):

You're welcome!

OpenStudy (xhawaiiangirlx):

Have a nice day!